A bakery finds that the price they can sell cakes is given by the function p = 580 − 10x where x is the number of cakes sold per day, and p is price. The total cost function of the company is given by c = (30+5x) 2 where x is previously defined, and c is total cost.
A. Find the revenue and marginal revenue functions [Hint: revenue is price multiplied by quantity i.e. revenue = price × quantity] (3 marks) B. Find the fixed cost and marginal cost function [Hint: fixed cost does not change with quantity produced] (3 marks) C. Find the profit function [Hint: profit is revenue minus total cost] (2 marks) D. Find the quantity that maximizes profit (2 marks)
1
Expert's answer
2020-05-05T20:32:33-0400

A.

The revenue function

"(580-10x)\cdot x=580x-x^2"

The marginal revenue function

"(580x-10x^2)'=580-2x"

B.

The fixed cost

"(30+5x)^2=900+300x+25x^2\\"

The fixed cost 900

The marginal cost function

"(900+300x+25x^2)'=300+50x"

C.

The profit function

"580x-10x^2-(30+5x)^2=-35x^2+280x-900\\"

D.

 The quantity that maximizes profit

"(-35x^2+280x-900)'=-70x+280\\\\-70x+280=0\\
x=4"